OpenStudy (anonymous):

EASY MEDAL solve 3x+5y=43 9x+7y=83

OpenStudy (anonymous):

@Compassionate @Donblue22 @shelbygt520 @sammixboo @shelbylynn @mathmath333 @math_1 @Theredman55555 @Secret-Ninja @tayromo @jordanloveangel @Tuta @fakeee @eliassaab

OpenStudy (jordanloveangel):

Solution for 9x+7=3x+43 equation: Simplifying 9x + 7 = 3x + 43 Reorder the terms: 7 + 9x = 3x + 43 Reorder the terms: 7 + 9x = 43 + 3x Solving 7 + 9x = 43 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. 7 + 9x + -3x = 43 + 3x + -3x Combine like terms: 9x + -3x = 6x 7 + 6x = 43 + 3x + -3x Combine like terms: 3x + -3x = 0 7 + 6x = 43 + 0 7 + 6x = 43 Add '-7' to each side of the equation. 7 + -7 + 6x = 43 + -7 Combine like terms: 7 + -7 = 0 0 + 6x = 43 + -7 6x = 43 + -7 Combine like terms: 43 + -7 = 36 6x = 36 Divide each side by '6'. x = 6 Simplifying x = 6

OpenStudy (jordanloveangel):

thx

OpenStudy (anonymous):

Am I missing something? Why are you solving @jordanloveangel 9x+7=3x+43

OpenStudy (mathmath333):

ur method is wrong @jordanloveangel

OpenStudy (jordanloveangel):

but its the right answer

OpenStudy (hockeychick23):

Ok so the process you want to use for this is elimination. Multiply the top equation by 3. 3(3x+5y=43) = 9x+15y=129 Now both equations have a 9x 9x+15y=129 9x+7y=83 Subtract the second equation from the first 8y=46 Divide both sides by 8. y= 5.75 To find x substitute y in x= 4.75

OpenStudy (anonymous):

I agree with @mathmath333

OpenStudy (jordanloveangel):

wait wait

OpenStudy (jordanloveangel):

Solution for 3x+5y=43 equation: Simplifying 3x + 5y = 43 Solving 3x + 5y = 43 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5y' to each side of the equation. 3x + 5y + -5y = 43 + -5y Combine like terms: 5y + -5y = 0 3x + 0 = 43 + -5y 3x = 43 + -5y Divide each side by '3'. x = 14.33333333 + -1.666666667y Simplifying x = 14.33333333 + -1.666666667y

OpenStudy (mathmath333):

9x + 7 \(\neq \) 3x + 43 \(43\neq 83\)

OpenStudy (jordanloveangel):

Solution for 9x+7y=83 equation: Simplifying 9x + 7y = 83 Solving 9x + 7y = 83 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-7y' to each side of the equation. 9x + 7y + -7y = 83 + -7y Combine like terms: 7y + -7y = 0 9x + 0 = 83 + -7y 9x = 83 + -7y Divide each side by '9'. x = 9.222222222 + -0.7777777778y Simplifying x = 9.222222222 + -0.7777777778y

OpenStudy (jordanloveangel):

okay @mathmath333 do the rest

OpenStudy (anonymous):

The solution by @hockeychick23 is right one.

OpenStudy (anonymous):

How can one get 2 medals for a solution that is not correct?

OpenStudy (hockeychick23):

The right answers are x= 4.75 and y= 5.75, plug in the answers and check